OpenStudy (anonymous):

Convert y=(3/4)x-5.5 into ax+by=d form.

OpenStudy (anonymous):

what math is this for

OpenStudy (anonymous):

IB math studies

OpenStudy (anonymous):

-3 x + 4 y = 22 ?

OpenStudy (anonymous):

Should be -22 on the left there.

OpenStudy (anonymous):

Well according to me the correct answer is 3x - 4y = 22

OpenStudy (anonymous):

yes 3x-4y=22 is the correct answer. this new equation doesn't make any sense. it doesn't give slope nor intercept. it gives nothing,so why you want to convert it in this form???

OpenStudy (anonymous):

y = (3/4) x - 5.5 Multiply each side by 4 4 y = 4*(3/4) x - 5.5*4 Evaluate the terms on the right side 4 y = (3) x - 22 Move (3)x to the left side -(3) x + 4 y = -(3) x + (3) x - 22 On the right side, -(3) x + (3) x cancel -3 x + 4 y = -22 Yes, polpak is mostly correct. The right side is -22. However, he said, "Should be -22 on the left there." I read polpak's comment at the time it was posted and ignored the entire comment content because of the left side error. Very sorry about all of this. I have put forth solutions on this site over 200 times and so far, this is the first time, that I know of, that I posted a solution with an error. I normally never attempt to solve problems posted on this site without the use of Mathematica to do the clerical work.

OpenStudy (anonymous):

Multiply both sides by -1 and you get 3x - 4y = 22, as I wrote.... We normaiil eliminate the minus if it occurs on both sides of the equation as in this case

OpenStudy (anonymous):

* normally

OpenStudy (anonymous):

Yes, you are correct about the result of the minus 1 multiplication. I guess everything conceivable about this problem has been laid to rest. Have a good day.

OpenStudy (anonymous):

Sorry, I'm good at math and bad at differentiating my right and left hands.